OpenStudy (anonymous):

Find the derivative dy/dx of x/y-x^2=1 by (a) solving each of the implicit equations for y explicitly in terms of x and (b) by differentiating each of the equations implicitly.

OpenStudy (anonymous):

so im not really sure exactly what the problem is asking but the implicit differentiation is (y-xy')/y^2 - 2x=0

OpenStudy (anonymous):

Thanks. :/ I have to solve it for y' thats what its asking :/

OpenStudy (anonymous):

oh thats easy. once you have (y-xy')/y^2 - 2x=0 all you do is add the 2x over then multiply everything by ysquared then substract the y over then divide by -x

OpenStudy (anonymous):

The answer is \[-2y ^{2}+ \frac{ y }{ x }\] If you could show me the steps how to get that it would solve the question for me.

OpenStudy (anonymous):

I solved it thanks to your steps! I cant believe it, I've been trying to solve this for 2h! :) Thanks a million!